OpenStudy (anonymous):

find the phase shift, period and amplitude of the function: y=1-2cos(2pix-3pi/2)

OpenStudy (mathmale):

A general formula for the cosine function is y = a*cos (bx + c). The AMPLITUDE of the function is the absolute value of ' a ': |a|. The PERIOD of the function is 2Pi/b, where b is the coefficient of the variable x. The PHASE SHIFT of the function is -c/b. The graph of y=1-2cos(2pix-3pi/2) involves a vertical translation, but you are not asked about that. Thus, you can safely focus on just y=-2cos(2pix-3pi/2).

OpenStudy (mathmale):

Using the formula I gave you, find the AMPLITUDE of this function. What is the coefficient of x in this function? Once you have that, find the PERIOD. What is ' c ' in this function? Once you have that, find the PHASE SHIFT.

OpenStudy (anonymous):

amplitude 2

OpenStudy (anonymous):

period is \[\frac{ 2\pi }{ 2 }\]

OpenStudy (anonymous):

phase shift is \[-\frac{ 7\pi }{ 2 }\]

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

But I see that the coefficient of your x is 2pi, not just 2. Would you mind calculating the period again?

OpenStudy (mathmale):

Also, I see that your ' c ' is -3pi/2. if your ' b ' is 2pi, What is your PHASE SHIFT, -c/b? I need to get off the Internet, but will be back on later today. I'll look for your responses. Take care! mathmale

OpenStudy (anonymous):

period is 1

OpenStudy (mathmale):

that's better. What about that phase shift?

OpenStudy (mathmale):

In summary: AMPLITUDE=2; PERIOD=1; PHASE SHIFT = ?

OpenStudy (anonymous):

\[-\frac{ 3 }{ 4 }\]

OpenStudy (mathmale):

Your ' c ' is -3Pi/2. What is -c? What is b? What is your PHASE SHIFT, -c/b?

OpenStudy (mathmale):

You were essentially right (about the PHASE SHIFT) the first time, but have a sign error there. Check that out. The phase shift in this problem is positive, not negative. sorry, I need to go. But again, if you respond, I'll respond back, later on today.

OpenStudy (anonymous):

3/4

OpenStudy (mathmale):

that's right! So, you've successfully answered this question. Awesome.